-Sport TANAPCALCBR1045.007 My Wise Rok Your teacher Need please answer question 3 and circle the answer please
Answer #1

Let the cutting sauase be ain Length = 11-za , width=11-2ae - height = a Length - 11-Qaso 1 llsea zasul az alg.g volume ofs! ਠੱਠ ਘm ਡ) =ਘਓ u: - ੧੫ ਓਪ ਨੂੰ = ਦ = ਪਰਖ ਠੇ = (ਹੋਠ- - - = ਪ2e ਠੇ = $6-)= - = ਪu ਗ ਹੋ = = 0 +ਹਉo s} OusmON = 0 07 (ਈN 0> hh

answered by: ANURANJAN SARSAM